# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

### v1

@@ VIII.2 Region 3: comparison with the analytical approximation
The first Figure VIII.5a, show $S_{,v}(v)$ and $\kappa_{-}(v)$ along the last-$u$ ray. They are indeed very close, as predicted by the analytical approximation. However, there’s a small visible gap near the end at around $v\sim 900$ (which is almost invisible in the corresponding Figure VII.9a). We will show in the next subsection that this is a finite-$\varepsilon_{0}$ deviation.
We see in the following Figure VIII.5b that the approximation is good for $S_{,u}$ as well. Similar to the $S_{,v}$case, there’s a small visible gap near the end at u=950u=-950 (which doesn’t appear at the corresponding Figure VII.9b). This is again a finite-$\varepsilon_{0}$ deviation.
Figure VIII.6: Analytical approximation of $R_{,v}$ and $R_{,u}$
@@
We move on to explore the analytical approximation for $R$. In Figure VIII.6a, we see that the analytical approximation for $R_{,v}$ Eq. (IV.40) works well. In this case, in contrast to Figure VII.10a, $R_{,v}>0$ because in the present case the chosen sources lead to $\widetilde{T}_{vv}<0$.
@@
Finally, we show in Figure VIII.6b that the analytical approximation for $R_{,u}$ Eq. (IV.41) works as well. This time, unlike Figure VII.10b (the case of vanishing sources), $\widetilde{T}_{uu}\neq 0$ because of the non-vanishing sources.
VIII.3 Impact of $\delta Q$
@@
In this run we used $\alpha=1$ parameter for $\delta Q$ (unlike the previous run which had $\alpha=0$). First we will show in Figure VIII.7 the value of $Q(u,v)$ along the last-$u$ ray (including the $\delta Q$ term) and compare it with $Q^{v}(v)$ which is the value of $Q$ at the EH. As seen in the figure, the value of $Q$ drifts by a lot and reaches 4 at maximal $v$, meaning that $Q(u,v)$ is greater than $M(v)$ – which might naively be regarded as “overcharging”.
Figure VIII.7: $Q^{v}$ and $Q=Q^{v}+\delta Q$ on the last u ray
@@
Recall that the analytical approximation for $S_{,v}$ is $\kappa_{-}^{v}$, and that $\kappa_{-}^{v}(v)$ was defined to be $\kappa_{-}^{v}(M(v),Q^{v}(v))$, see Eq. (IV.6). One might be tempted to think that perhaps the more relevant charge parameter, to be used in $\kappa_{-}^{v}(v)$, would be $Q(u,v)=Q^{v}(v)+\delta Q(u,v)$ rather than just $Q^{v}(v)$. The purpose of the next figure is to check this issue.
@@
In Figure VIII.8 we show $S_{,v}$ along with two “candidate” $\kappa_{-}^{v}$ curves: the green with the correct choice $Q^{v}(v)$, and red with the incorrect choice $Q(u,v)=\delta Q+Q^{v}$. As can be clearly seen in the figure, the green curve matches the blue curve, while the red curve quickly deviates from them in region 3. (The red curve is cut short because $M$ becomes bigger than $Q(u,v)$, meaning the corresponding $\kappa_{-}$ would have a negative square root.) This figure indicates that using $Q^{v}(v)$ in the analytical approximation is indeed the correct choice. This is not surprising, of course, because our construction of the analytical approximation for $S_{,v}$ in subsection IV.3 provided the $S_{,v}=-\kappa_{-}^{v}$ specifically with the charge parameter $Q^{v}(v)$, see Eq. (IV.6).
Figure VIII.8: $S_{,v}$ and $\kappa_{-}$ on last u ray, using both $Q^{v}$ and $Q=\delta Q+Q^{v}$ for $\kappa_{-}$
VIII.4 Deviations from the analytical approximation
@@
In this section we intend to analyze the deviation from the analytical approximation in order to test whether it is a finite-$\varepsilon_{0}$ effect. To this end, similar to subsection VII.4 above, we’ll compare the deviation from the analytical approximation for two different $\varepsilon_{0}$ values: $\frac{1}{500}$ and $\frac{1}{1000}$. We will again scale the $\varepsilon_{0}=\frac{1}{500}$ curve by a factor of $\frac{1}{2}$ for $S$ and $\frac{1}{4}$ for $R$. As before, we will plot these scaled deviations as a function of the mass $M(v)$. We will see that the two curves overlap, confirming that the deviation scales as $\varepsilon_{0}$ for $S$ and as $\varepsilon_{0}^{2}$ for $R$. We shall also show the unscaled $\varepsilon_{0}=\frac{1}{500}$ curve in orange.
Figure VIII.9: $S_{,v}+\kappa_{-}$ with two $\varepsilon_{0}$, scaled by $\frac{1}{2}$ on last u ray
@@
The first Figure VIII.9 shows the deviation of $S_{,v}$ for the two mentioned values of $\varepsilon_{0}$, with the $\varepsilon_{0}=\frac{1}{500}$ deviations scaled by $\frac{1}{2}$. The two curves (green and blue) indeed overlap, confirming the scaling of the deviation as $\varepsilon_{0}.$ Comparing to Figure VII.13 (vanishing sources run), we see that the deviation is larger by around a factor of $\sim 2$ in the current case, but this change is of no importance.
Figure VIII.10: $S_{,u}+\kappa_{-}^{u}$ with two $\varepsilon_{0}$, scaled by $\frac{1}{2}$ on last v ray
@@
The second Figure VIII.10 shows the deviation of $S_{,u}$ for the same two values of $\varepsilon_{0}$, with the $\varepsilon_{0}=\frac{1}{500}$ deviations scaled by $\frac{1}{2}$. The two curves (blue and green) again overlap, confirming a scaling of the deviation in $S_{,u}$ as $\varepsilon_{0}$.
@@
Figure VIII.11: $R_{,v}+\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ with two $\varepsilon_{0}$ and scaled by $\frac{1}{4}$ on last u ray
@@
The following Figure VIII.11 shows the deviation of $R_{,v}$ for the two values of $\varepsilon_{0}$, with $\varepsilon_{0}=\frac{1}{500}$ curve scaled by $\frac{1}{4}$. The two curves overlap, confirming scaling with $\varepsilon_{0}^{2}$ – as expected from $R$ derivatives in region 3, see explanation in subsection VII.4.
@@
Finally we explore the deviation of $R_{,u}$ from its analytical approximation. Unlike the no-sources run considered in subsection VII.4, this time there’s non-zero $\widetilde{T}_{uu}$ and hence non-vanishing $R_{,u}$ in region 3, and Figure VIII.12 shows the deviation of $R_{,u}$, with the $\varepsilon_{0}=\frac{1}{500}$ curve scaled by $\frac{1}{4}$. The two curves (green and blue) again overlap, confirming the scaling of the deviation as $\text{$\varepsilon_{0}^{2}$}.$ (We expect the same scaling of the deviation for $R_{,u}$ as for $R_{,v}$.)
@@
Figure VIII.12: $R_{,u}+\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ with two $\varepsilon_{0}$ and scaled by $\frac{1}{4}$ on last v ray
IX Finite-$du$ and finite-$T_{0}$ errors
@@
Specifically, we will work here with the set of parameters of section VIII:
@@
IX.1 Finite-$du$ error
Our algorithm discretizes spacetime into a two-dimensional grid with separation $du$ (we use $du=dv$) between two adjacent grid locations. Because the distance between grid points is finite, we have an error in our numerical calculation, which increases with the size of $du$ – the so called truncation error. The smaller $du$ is, the smaller the truncation error gets. We would like to estimate this error which results from having finite $du$.
To estimate the magnitude of this error, we run the numerical code with three different $du$ values: $du=\frac{1}{50}$, $du=\frac{1}{100}$, and $du=\frac{1}{200}$, and compare the results. The difference between the runs $du=\frac{1}{50}$ and $du=\frac{1}{100}$ provides an estimate for the truncation error in the $du=\frac{1}{50}$ run, and similarly the difference between the $du=\frac{1}{100}$ and $du=\frac{1}{200}$ runs provides an estimate for the truncation error in the $du=\frac{1}{100}$ run, as we will now explain.
Let $f\left(du\right)$ denote some numerically computed quantity (that necessarily depends on $du$), and let $e\left(du\right)$ denote the numerical error in the computation of that quantity. We expect the error to scale as $du^{2}$, as already mentioned above in section V (and as will be confirmed below):
$e\left(du\right)=\beta\cdot du^{2}$ (IX.1)
@@ IX.1 Finite-$du$ error
$f\left(du\right)-f\left(\frac{1}{2}du\right)$ $=e\left(du\right)-e\left(\frac{1}{2}du\right)$ (IX.2)
$=\beta\left(du^{2}-\left(\frac{1}{2}du\right)^{2}\right)=\frac{3}{4}\beta du^{2}$ (IX.3)
@@
$\beta=\frac{4}{3}\frac{f\left(du\right)-f\left(\frac{1}{2}du\right)}{du^{2}}$ (IX.4)
@@
$e\left(du\right)=\frac{4}{3}\left(f\left(du\right)-f\left(\frac{1}{2}du\right)\right)$ (IX.5)
@@
$e\left(\frac{1}{100}\right)=\frac{4}{3}\left[f\left(\frac{1}{100}\right)-f\left(\frac{1}{200}\right)\right]$ (IX.6)
@@
Figure IX.1: Truncation error in $R$ on last u ray
We first present the results for the truncation error in $R$ in Figure IX.1, which shows the two differences between running with different $du$ values. Note how in region 1 the error is very close to zero. The error for $du=\frac{1}{100}$ is of order $\sim 10^{-5}$ as indicated by the orange curve.
Figure IX.2: Ratio of truncation errors $\frac{R\left[du=\frac{1}{50}\right]-R\left[du=\frac{1}{100}\right]}{R\left[du=\frac{1}{100}\right]-R\left[du=\frac{1}{200}\right]}$ on last u ray
We next present the ratio of these two differences in Figure IX.2. As expected, the ratio is equal to 4, confirming error scaling as $du^{2}$. We point out that the error in region 1 (not seen in this figure) is exceptional: it is very small (see Figure IX.1) and it doesn’t scale as $du^{2}$. In this region, the truncation error is negligible and other errors seem to dominate.
@@
Figure IX.3: Truncation error in S on last u ray
Figure IX.3 shows the truncation error in $S$. Note that while the error reaches a relatively high value, 0.04, $S$ itself reaches $\sim 4000$ so the relative error is again of order $10^{-5}$.
Figure IX.4: Ratio of truncation errors $\frac{S\left[du=\frac{1}{50}\right]-S\left[du=\frac{1}{100}\right]}{S\left[du=\frac{1}{100}\right]-S\left[du=\frac{1}{200}\right]}$ on last u ray
We display the ratio of these two truncation errors in Figure IX.4, and it can be seen that in region 2 and 3 the ratio is 4, confirming error scaling with $du^{2},$as before. In region 1, the error is very small and doesn’t scale with $du$ for the same reason explained above.
Figure IX.5: Truncation error in $R_{,v}$ on last u ray
Next we show the errors in $R_{,v}$ in Figure IX.5. In region 1, the errors are roughly equal because truncation error isn’t the leading cause of errors. In region 3 they seem to scale as expected. We computed the ratio in Figure IX.6, and it appears concentrated around 4, as expected from $du^{2}$ error scaling. It appears more jittery because of round-off error.
Figure IX.6: Ratio of truncation errors $\frac{R_{,v}\left[du=\frac{1}{50}\right]-R_{,v}\left[du=\frac{1}{100}\right]}{R_{,v}\left[du=\frac{1}{100}\right]-R_{,v}\left[du=\frac{1}{200}\right]}$ on last u ray
Figure IX.7: Truncation error in $S_{,v}$ on last u ray
In Figure IX.7 we show the error in $S_{,v}$. It is of order $10^{-4}.$ The error in region 1 is again close to zero.
Figure IX.8: Ratio of truncation errors $\frac{S_{,v}\left[du=\frac{1}{50}\right]-S_{,v}\left[du=\frac{1}{100}\right]}{S_{,v}\left[du=\frac{1}{100}\right]-S_{,v}\left[du=\frac{1}{200}\right]}$ on last u ray
We see in Figure IX.8 that the ratio is also around 4 in region 3, as expected from $du^{2}$ scaling.

### v2

@@ 8.2 Region 3: comparison with the analytical approximation
The first Figure 8.5a, show $S_{,v}(v)$ and $\kappa_{-}(v)$ along the last-$u$ ray. They are indeed very close, as predicted by the analytical approximation. However, there is a small visible gap near the end at around $v\sim 900$ (which is almost invisible in the corresponding Figure 7.9a). We will show in the next subsection that this is a finite-$\varepsilon_{0}$ deviation.
We see in the following Figure 8.5b that the approximation is good for $S_{,u}$ as well. Similar to the $S_{,v}$case, there is a small visible gap near the end at u=950u=-950 (which doesn’t appear at the corresponding Figure 7.9b). This is again a finite-$\varepsilon_{0}$ deviation.
Figure 8.6: Analytical approximation of $R_{,v}$ and $R_{,u}$
@@
We move on to explore the analytical approximation for $R$. In Figure 8.6a, we see that the analytical approximation for $R_{,v}$ Eq. (4.40) works well. In this case, in contrast to Figure 7.10a, $R_{,v}>0$ because in the present case the chosen sources lead to $\widetilde{T}_{vv}<0$.
@@
Finally, we show in Figure 8.6b that the analytical approximation for $R_{,u}$ Eq. (4.41) works as well. This time, unlike Figure 7.10b (the case of vanishing sources), $\widetilde{T}_{uu}\neq 0$ because of the non-vanishing sources.
8.3 Impact of $\delta Q$
@@
In this run we used $\alpha=1$ parameter for $\delta Q$ (unlike the previous run which had $\alpha=0$). First we will show in Figure 8.7 the value of $Q(u,v)$ along the last-$u$ ray (including the $\delta Q$ term) and compare it with $Q^{v}(v)$ which is the value of $Q$ at the EH. As seen in the figure, the value of $Q$ drifts by a lot and reaches 4 at maximal $v$, meaning that $Q(u,v)$ is greater than $M(v)$ – which might naively be regarded as “overcharging”.
Figure 8.7: $Q^{v}$ and $Q=Q^{v}+\delta Q$ on the last u ray
@@
Recall that the analytical approximation for $S_{,v}$ is $\kappa_{-}^{v}$, and that $\kappa_{-}^{v}(v)$ was defined to be $\kappa_{-}^{v}(M(v),Q^{v}(v))$, see Eq. (4.6). One might be tempted to think that perhaps the more relevant charge parameter, to be used in $\kappa_{-}^{v}(v)$, would be $Q(u,v)=Q^{v}(v)+\delta Q(u,v)$ rather than just $Q^{v}(v)$. The purpose of the next figure is to check this issue.
@@
In Figure 8.8 we show $S_{,v}$ along with two “candidate” $\kappa_{-}^{v}$ curves: the green with the correct choice $Q^{v}(v)$, and red with the incorrect choice $Q(u,v)=\delta Q+Q^{v}$. As can be clearly seen in the figure, the green curve matches the blue curve, while the red curve quickly deviates from them in region 3. (The red curve is cut short because $M$ becomes bigger than $Q(u,v)$, meaning the corresponding $\kappa_{-}$ would have a negative square root.) This figure indicates that using $Q^{v}(v)$ in the analytical approximation is indeed the correct choice. This is not surprising, of course, because our construction of the analytical approximation for $S_{,v}$ in subsection 4.3 provided the $S_{,v}=-\kappa_{-}^{v}$ specifically with the charge parameter $Q^{v}(v)$, see Eq. (4.6).
Figure 8.8: $S_{,v}$ and $\kappa_{-}$ on last u ray, using both $Q^{v}$ and $Q=\delta Q+Q^{v}$ for $\kappa_{-}$
8.4 Deviations from the analytical approximation
@@
In this section we intend to analyze the deviation from the analytical approximation in order to test whether it is a finite-$\varepsilon_{0}$ effect. To this end, similar to subsection 7.4 above, we’ll compare the deviation from the analytical approximation for two different $\varepsilon_{0}$ values: $\frac{1}{500}$ and $\frac{1}{1000}$. We will again scale the $\varepsilon_{0}=\frac{1}{500}$ curve by a factor of $\frac{1}{2}$ for $S$ and $\frac{1}{4}$ for $R$. As before, we will plot these scaled deviations as a function of the mass $M(v)$. We will see that the two curves overlap, confirming that the deviation scales as $\varepsilon_{0}$ for $S$ and as $\varepsilon_{0}^{2}$ for $R$. We shall also show the unscaled $\varepsilon_{0}=\frac{1}{500}$ curve in orange.
Figure 8.9: $S_{,v}+\kappa_{-}$ with two $\varepsilon_{0}$, scaled by $\frac{1}{2}$ on last u ray
@@
The first Figure 8.9 shows the deviation of $S_{,v}$ for the two mentioned values of $\varepsilon_{0}$, with the $\varepsilon_{0}=\frac{1}{500}$ deviations scaled by $\frac{1}{2}$. The two curves (green and blue) indeed overlap, confirming the scaling of the deviation as $\varepsilon_{0}.$ Comparing to Figure 7.13 (vanishing sources run), we see that the deviation is larger by around a factor of $\sim 2$ in the current case, but this change is of no importance.
Figure 8.10: $S_{,u}+\kappa_{-}^{u}$ with two $\varepsilon_{0}$, scaled by $\frac{1}{2}$ on last v ray
@@
The second Figure 8.10 shows the deviation of $S_{,u}$ for the same two values of $\varepsilon_{0}$, with the $\varepsilon_{0}=\frac{1}{500}$ deviations scaled by $\frac{1}{2}$. The two curves (blue and green) again overlap, confirming a scaling of the deviation in $S_{,u}$ as $\varepsilon_{0}$.
@@
Figure 8.11: $R_{,v}+\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ with two $\varepsilon_{0}$ and scaled by $\frac{1}{4}$ on last u ray
@@
The following Figure 8.11 shows the deviation of $R_{,v}$ for the two values of $\varepsilon_{0}$, with $\varepsilon_{0}=\frac{1}{500}$ curve scaled by $\frac{1}{4}$. The two curves overlap, confirming scaling with $\varepsilon_{0}^{2}$ – as expected from $R$ derivatives in region 3, see explanation in subsection 7.4.
@@
Finally we explore the deviation of $R_{,u}$ from its analytical approximation. Unlike the no-sources run considered in subsection 7.4, this time there is non-zero $\widetilde{T}_{uu}$ and hence non-vanishing $R_{,u}$ in region 3, and Figure 8.12 shows the deviation of $R_{,u}$, with the $\varepsilon_{0}=\frac{1}{500}$ curve scaled by $\frac{1}{4}$. The two curves (green and blue) again overlap, confirming the scaling of the deviation as $\text{$\varepsilon_{0}^{2}$}.$ (We expect the same scaling of the deviation for $R_{,u}$ as for $R_{,v}$.)
@@
Figure 8.12: $R_{,u}+\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ with two $\varepsilon_{0}$ and scaled by $\frac{1}{4}$ on last v ray
9 Finite-$du$ and finite-$T_{0}$ errors
@@
Specifically, we will work here with the set of parameters of section 8:
@@
9.1 Finite-$du$ error
Our algorithm discretizes spacetime into a two-dimensional grid with separation $du$ (we use $du=dv$) between two adjacent grid locations. Because the distance between grid points is finite, we have an error in our numerical calculation, which increases with the size of $du$. We would like to estimate this error which results from having finite $du$.
To estimate the magnitude of this error, we run the numerical code with three different $du$ values: $du=\frac{1}{50}$, $du=\frac{1}{100}$, and $du=\frac{1}{200}$, and compare the results. The difference between the runs $du=\frac{1}{50}$ and $du=\frac{1}{100}$ provides an estimate for the truncation error in the $du=\frac{1}{50}$ run (up to a factor $\frac{4}{3}$ which we explain below), and similarly the difference between the $du=\frac{1}{100}$ and $du=\frac{1}{200}$ runs provides an estimate for the truncation error in the $du=\frac{1}{100}$ run, as we will now explain.
Let $f\left(du\right)$ denote some numerically computed quantity (that necessarily depends on $du$), and let $e\left(du\right)$ denote the numerical error in the computation of that quantity. We expect the error to scale as $du^{2}$, as already mentioned above in section 5 (and as will be confirmed below):
$e\left(du\right)=\beta\cdot du^{2}$ (9.1)
@@ 9.1 Finite-$du$ error
$f\left(du\right)-f\left(\frac{1}{2}du\right)$ $=e\left(du\right)-e\left(\frac{1}{2}du\right)$ (9.2)
$=\beta\left(du^{2}-\left(\frac{1}{2}du\right)^{2}\right)=\frac{3}{4}\beta du^{2}$ (9.3)
@@
$\beta=\frac{4}{3}\frac{f\left(du\right)-f\left(\frac{1}{2}du\right)}{du^{2}}$ (9.4)
@@
$e\left(du\right)=\frac{4}{3}\left(f\left(du\right)-f\left(\frac{1}{2}du\right)\right)$ (9.5)
@@
$e\left(\frac{1}{100}\right)=\frac{4}{3}\left[f\left(\frac{1}{100}\right)-f\left(\frac{1}{200}\right)\right]$ (9.6)
@@
Figure 9.1: Truncation error in $R$ on last u ray
We first present the results for the truncation error in $R$ in Figure 9.1, which shows the two differences between running with different $du$ values. Note how in region 1 the error is very close to zero. The error for $du=\frac{1}{100}$ is of order $\sim 10^{-5}$ as indicated by the orange curve.
Figure 9.2: Ratio of truncation errors $\frac{R\left[du=\frac{1}{50}\right]-R\left[du=\frac{1}{100}\right]}{R\left[du=\frac{1}{100}\right]-R\left[du=\frac{1}{200}\right]}$ on last u ray
We next present the ratio of these two differences in Figure 9.2. As expected, the ratio is equal to 4, confirming error scaling as $du^{2}$. We point out that the error in region 1 (not seen in this figure) is exceptional: it is very small (see Figure 9.1) and it doesn’t scale as $du^{2}$. In this region, the truncation error is negligible and other errors seem to dominate.
@@
Figure 9.3: Truncation error in S on last u ray
Figure 9.3 shows the truncation error in $S$. Note that while the error reaches a relatively high value, 0.04, $S$ itself reaches $\sim 4000$ so the relative error is again of order $10^{-5}$.
Figure 9.4: Ratio of truncation errors $\frac{S\left[du=\frac{1}{50}\right]-S\left[du=\frac{1}{100}\right]}{S\left[du=\frac{1}{100}\right]-S\left[du=\frac{1}{200}\right]}$ on last u ray
We display the ratio of these two truncation errors in Figure 9.4, and it can be seen that in region 2 and 3 the ratio is 4, confirming error scaling with $du^{2},$as before. In region 1, the error is very small and doesn’t scale with $du$ for the same reason explained above.
Figure 9.5: Truncation error in $R_{,v}$ on last u ray
Next we show the errors in $R_{,v}$ in Figure 9.5. In region 1, the errors are roughly equal because truncation error isn’t the leading cause of errors. In region 3 they seem to scale as expected. We computed the ratio in Figure 9.6, and it appears concentrated around 4, as expected from $du^{2}$ error scaling. It appears more jittery because of round-off error.
Figure 9.6: Ratio of truncation errors $\frac{R_{,v}\left[du=\frac{1}{50}\right]-R_{,v}\left[du=\frac{1}{100}\right]}{R_{,v}\left[du=\frac{1}{100}\right]-R_{,v}\left[du=\frac{1}{200}\right]}$ on last u ray
Figure 9.7: Truncation error in $S_{,v}$ on last u ray
In Figure 9.7 we show the error in $S_{,v}$. It is of order $10^{-4}.$ The error in region 1 is again close to zero.
Figure 9.8: Ratio of truncation errors $\frac{S_{,v}\left[du=\frac{1}{50}\right]-S_{,v}\left[du=\frac{1}{100}\right]}{S_{,v}\left[du=\frac{1}{100}\right]-S_{,v}\left[du=\frac{1}{200}\right]}$ on last u ray
We see in Figure 9.8 that the ratio is also around 4 in region 3, as expected from $du^{2}$ scaling.